OpenStudy (anonymous):

I WILL FAN AND MEDAL! Find the product of (3x + 7y)(3x − 7y). 9x2 − 42xy + 49y2 9x2 + 42xy + 49y2 9x2 − 49y2 9x2 + 49y2

OpenStudy (anonymous):

cant go farther than that

OpenStudy (anonymous):

Do you know foil?

OpenStudy (anonymous):

Get rid of the squares

OpenStudy (anonymous):

cant solve it any more

OpenStudy (anonymous):

\[\large (a+b)(a-b)=a^2-b^2\] with \[a=3x,b=7y\]

OpenStudy (anonymous):

You cant get rid of them bcus they have to be the same variables

OpenStudy (anonymous):

All of the 2's are exponets and @domu i do not know foil

OpenStudy (anonymous):

\[(x+a)(x-a) = x^2 - a^2 \]

OpenStudy (anonymous):

as usual it is C because it is always C every time it is C never A or B once in a while D well, not too often

OpenStudy (anonymous):

I'm sorry guys im not understanding where the a's and b's are coming from :(

OpenStudy (anonymous):

[9x^{2}-49y ^{2}\] is correct by using formula

OpenStudy (anonymous):

(3x + 7y)(3x − 7y) 3x * 3x = 9x^2 3x * -7y = -21xy 7y * 3x = +21xy 7y * -7y = -49y 9x^2 -21xy + 21xy - 49y 9x^2 - 49y

OpenStudy (anonymous):

@SOS101

OpenStudy (anonymous):

Thank You all!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

did i mention is was C?

OpenStudy (anonymous):

LOL @satellite73 yes u did Sir

OpenStudy (anonymous):

so long method of @some_someone

OpenStudy (anonymous):

the \(a\) and \(b\) are general it is always the case that \[(a+b)(a-b)=a^2-b^2\] and so \[(3x+7y)(3x-7y)=(3x)^2-(7y)^2=9x^2-49y^2\]

OpenStudy (anonymous):

Lol im sorry satellite I was trying to give the person with less points a medal, Sorry for the inconvience im laughing so hard at satellite's comment

OpenStudy (anonymous):

I will get my sister to give u a medal @satellite73

OpenStudy (anonymous):

7y*7y=49y^2?????

OpenStudy (anonymous):

i have plenty, thank you though and my sincere apologies to all who found my post abusive ...

OpenStudy (anonymous):

@rashad.dawkins do you have a question?

OpenStudy (anonymous):

Lol No @satellite73 Not abusive ; just really had me crackin up

OpenStudy (anonymous):

I thought 7y^2 times itself is 49 squared?